ABC is an isosceles triangle and AC,BC are the tangents at M and N respectively.DE is the diameter of the circle.Angle ADP=AngleBEQ=100 degree.what is the value of angle prd

Grade:12th pass

1 Answers

Vetsa Sreekar
4 years ago
Given that angle ADP = angle BEQ = 100
angle PDE = angle QED = 180 - 100 = 80 (linear pair)
∆PDE and ∆QED are right angled triangles (angle in a semi circle is 90)
angle EDQ = angle DEP = 10 ( angle sum property)
In ∆RDE, angle DRB = 180 - 10 -10 = 160 ( angle sum property)
anglePRD = 180 - 160 = 20 (linear pair)​